Campus email text:
Saturday, January 10, Technology Services will be performing important updates on our database systems. There will be two groups of outages:
5 AM until 8 PM – Cascade and Cascade Web will be unavailable.
9 AM until Noon – Banner, Famis, Basis, and Millennium will be unavailable.
Email, network files and shares, and other network services will not be impacted.
Cascade database is switching to new hardware and upgrading to 10g. Administrative toolset is being upgraded to 10g and deployed on new hardware.
Banner, Famis, Basis, and Millennium will be patched to make them compatible with 10g.
All Sophos pure message anti-spam and anti-virus email servers will be down for maintenance from 5:30-7:30am to upgrade the software to the latest version. This upgrade is required to improve spam blocking for UPS email users.
During the upgrade mail should continue to flow between the University and the Internet, but may be slightly delayed at times until the upgrade is complete at 7:30am. The spam quarantine will be unavailable during parts of this outage as well. If the quarantine is not available please check back later in the day.
Update — 10/14/2008 @ 6:54am
All upgrades are complete and working as expected. Due to the upgrade process some email may have been delayed by a half hour or more, but should now be back to normal.
Starting at 6:30am on Friday, October 10th, the Sophos Spam Quarantine will be down for maintenance. During this time users will not be able to access their spam quarantines, but mail will continue to be delivered as normal. The Quarantine is expected to be available by 8:30am.
Update — 10/10/2008 @ 8:30am
Migration is taking longer than expected waiting on replication of small files. We currently estimate that the Spam Quarantine will be up approximately 9:30am.
Update — 10/10/2008 @ 9:45am
Migration was completed at approximately 9:30am.
Next week we will be moving web content on www2.ups.edu to new web server hardware. This is a routine replacement of an older web server. The move will take place on Tuesday evening, August 19, beginning at 5 PM and ending early morning on the next day, August 20. During this time it will not be possible to update websites on this server; however, the website itself will be available for normal viewing at all times. This will have no effect on content served from the main University web server, www.ups.edu.
We will be making a change to our Internet router on 08/06/2008 between 6:00-7:00am. During this time you would experience intermittent outages to the internet. Please be advised that connectivity cannot be guaranteed during this time. Thank you for your understanding.
We will be replacing the main firewall on 07/31/2008 between 6-7:00am. There will be intermittent outages as we shutdown the old firewall and bring up the new one.
I will be upgrading the network firewall device on 07/24/2008 at 6:00-7:00am. Expect intermittent internet outages during this time.
We have successfully upgraded our network Internet Border Router. As you might remember we were experiencing problems with our Internet connection due to our existing Border Router being undersized. We have now successfully upgraded the hardware to properly handle the internet traffic we have experienced and what is projected for this comming academic year. We hope you will realize higher reliability with this new equipment in place.
We replaced the network switches in the south end of Wyatt Hall this week-end. The north end (classrooms) will be replaced this week (during spring break).
Blackboard will be fitted with some additional disk space to provide for more courses and content. To ensure a smooth upgrade, we have allotted 2 days for the procedure. Blackboard will be unavailable starting at 8:00 AM Thursday the 3rd, and return to service no later than 5:00 PM on Friday, the 4th.
Wireless coverage has been expanded to cover all of Howarth Hall.
The movement of files from the old Merlin2 server was completed successfully. The new server is ready and available.
It was necessary to replace the SCSI backplane in the web server in order to bring one of the disk arrays back online. A service technician replaced the hardware, and the array was rebuilt.
DBA reported problems with CRYSTAL on Saturday, 8/28 in the evening.
Examined system on Monday 8/30 system appears to be unable to communicate with SAN. Cleaned fibre, switch and HBA but no luck.
Examined HBA on 8/31 no LED on card. Called vendor support–3.5 hours later the HBA was considered bad and a new HBA was sent out–4 hour delivery. New HBA installed, but still loading SAN drives. Lights are now working on both HBA and switch.
Called back support engineer at 9am left voice message. Called vendor support at 11 am to talk with another engineer and was told our call would be assigned to another engineer. Original engineer called back at 4pm to apologize that another engineer had not been assigned. Dual entries seen in switch, old entry removed and system rebooted and SAN drives are not visible.
A hard disk failure in the RAID array on the media server resulted in the server hanging. After making several attempts to recover the system, the data was backed up and system was rebuilt.